# Organizing Unstructured Image Collections using Natural Language

### v1

@@ E.3 Supplementary Results for Comparison Criteria-Conditioned Clustering Methods
Appendix F Further Studies of the Proposed Framework
In this section, we provide additional studies on our proposed framework, using the main configuration (Caption-based Proposer and Caption-based Grouper) for the analysis. In Sec. F.1, we conduct control experiments to examine the system sensitivity of our framework to different multimodal large language models (MLLMs) and large language models (LLMs). In Sec. F.2, we demonstrate how incorporating advanced prompting strategies can further enhance the framework’s performance on fine-grained criteria.
F.1 Further System Sensitivity Analysis of Various MLLMs and LLMs
@@
Figure 10: Sensitivity analysis of different MLLMs and LLMs on the six SMC benchmarks. Top (a): We fix the LLM to Llama-3.1-8B and study the impact of different MLLMs. Bottom (b): We fix the MLLM to LLaVA-NeXT-7B and study the impact of different LLMs. The average clustering accuracy(%) across different criteria is reported on the left, while the average semantic accuracy(%) is reported on the right.
@@
In Fig. 10, we perform a system-level sensitivity analysis using our default system configuration (caption-based proposer and caption-based grouper) to examine the impact of different MLLMs and LLMs on the system performance. Since all variants successfully propose the basic criteria in each benchmark, we report the average clustering accuracy (CAcc) and semantic accuracy (SAcc) across various criteria for comparative analysis.
@@
Specifically, in Fig. 10(a), we first fix the LLM in our system to Llama-3.1-8B (Meta, 2024b) and assess the influence of various MLLMs: GPT-4V (OpenAI, 2023), BLIP-3-4B (Xue et al., 2024), and LLaVA-NeXT-7B (Liu et al., 2024b). Next, in Fig. 10(b), we set the MLLM to LLaVA-NeXT-7B and evaluate different LLMs: GPT-4-turbo (OpenAI, 2023), GPT-4o (OpenAI, 2024), Llama-3-8B (Meta, 2024a), and Llama-3.1-8B.
@@
Findings in Fig. 10(a) indicate a direct correlation between the size of the MLLM and the ability of our system to uncover substructures, highlighting the significant role of MLLMs in translating visual information into natural language. On the other hand, this scalability demonstrates that our system can enhance performance with more robust MLLMs, thanks to its training-free design, which ensures compatibility with any MLLM. Despite this, we use LLaVA-NeXT-7B as our default MLLM due to its reproducibility, being open-source and unaffected by API changes, and its capacity for local deployment, which upholds privacy by not exposing sensitive image data to external entities.
@@
As for the LLMs, as depicted in Fig. 10(b), despite GPT-4-turbo showing marginally superior performance, the open-source Llama-3.1-8B achieves similar results across benchmarks, making it our default LLM. Notably, except for the Card-2c dataset, system performance remains largely consistent regardless of the power of the LLM. This consistency suggests that the reasoning task for SMC, given the capabilities of modern LLMs to tackle complex problems (Street et al., 2024), is relatively straightforward.
F.2 Further Study on Fine-grained Image Collections
@@
Image collections may include fine-grained grouping criteria, such as Bird species in bird photography. Fine-grained criteria pose unique challenges for substructure discovery due to small inter-class differences and large intra-class variations (Zhang et al., 2014; Vedaldi et al., 2014; He & Peng, 2017). This requires the model to detect subtle visual distinctions to accurately infer cluster names and guide the grouping process. The straightforward captioning process in our current framework may not fully capture these subtle visual nuances. However, the modular design of our framework allows for seamless integration of advanced cross-modal chain-of-thought (CoT) prompting strategies to address this issue.
@@
Appendix G Further Qualitative Results
In this section, we visualize the grouping results predicted by the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper). Specifically, we present example clustering results across different criteria for COCO-4c in Fig. 11, Food-4c in Fig. 12, Action-3c in Fig. 13, Clevr-4c in Fig. 14, and Card-2c in Fig. 15. Additionally, we showcase example clustering results at different predicted granularity levels for COCO-4c in Fig. 16.
Figure 11: Example predicted clusters of COCO-4c.
Figure 12: Example predicted clusters of Food-4c.
Figure 13: Example predicted clusters of Action-3c.
Figure 14: Example predicted clusters of Clevr-4c.
Figure 15: Example predicted clusters of Card-2c.
Appendix H Further Failure Case Analysis
In Fig. 17, we present several failure cases from the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper). As observed, our method frequently mis-assigns “Surfing” to the “Kayaking” cluster under the Activity criterion. Upon examining the intermediate criterion captions generated by the MLLM, we found that this error is largely due to the MLLM incorrectly describing a “Surfboard” as a “Kayak”. This highlights the importance of the MLLM’s ability to accurately describe images, as it is critical for the performance of our system. Potential improvements could include majority voting or model ensembling using different MLLM models.
Finally, we observed that our method sometimes fails to distinguish subtle, fine-grained differences between images, leading to incorrect labels. For example, as shown in Fig. 17, “Edamame” or “Pho” are typical dishes from China, Vietnam, and Japan, but they may be presented differently depending on the cuisine. The “Edamame” shown in Fig. 17 is presented in a traditional Japanese style, yet our model incorrectly predicted it as Chinese cuisine. This oversight of fine-grained details could be improved by employing a more advanced prompting strategy (Liu et al., 2024e).
Appendix I Further Details on Application Study
@@
In this section, we present additional implementation details, evaluation results, and findings for the application study discussed in Sec. 5 of the main paper. Specifically, Sec. I.1 offers further evaluation results and implementation details on using our predicted distribution to train a debiased model with GroupDRO (Sagawa et al., 2020). Sec. I.2 outlines the implementation of the user study that assesses the alignment between predicted biases and human judgments, along with comprehensive findings for all studied occupations and identified criteria. Finally, Sec. I.3 provides additional insights from the analysis of social media image popularity.
I.1 Further Details on Discovering and Mitigating Dataset Bias
In this section, we provide additional evaluation results and implementation details for the application study presented in Sec. 5 of the main paper.
@@ I.1 Further Details on Discovering and Mitigating Dataset Bias
where $H(A)$ and $H(Y)$ represent the normalized entropy of the gender and hair color distributions, respectively. A value of $H(A)$ or $H(Y)$ equal to 1 indicates a uniform distribution (i.e., no class imbalance). We then used the ground-truth distribution from the dataset’s labels and our predicted distribution to estimate the spurious correlation intensity using the score from Eq. 1. For gender and hair color, our method’s predictions yielded a score of $I_{Pred}=0.10$, which is nearly identical to the ground-truth score of $I_{GT}=0.11$. This demonstrates that our method effectively identifies and confirms the bias directly from the training set.

### v4

@@ G.3 Further Comparative Results with TCMC Methods
Appendix H Qualitative Analysis
In this section, we visualize the grouping results predicted by the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper). Specifically, we present example clustering results across different criteria for COCO-4c in Figure 10, Food-4c in Figure 11, Action-3c in Figure 12, Clevr-4c in Figure 13, and Card-2c in Figure 14. Additionally, we showcase example clustering results at different predicted granularity levels for COCO-4c in Figure 15.
Figure 10: Example predicted clusters of COCO-4c.
Figure 11: Example predicted clusters of Food-4c.
Figure 12: Example predicted clusters of Action-3c.
Figure 13: Example predicted clusters of Clevr-4c.
Figure 14: Example predicted clusters of Card-2c.
Appendix I Failure Case Analysis
In Figure 16, we present several failure cases from the best configuration of our proposed framework (Caption-based Proposer and Caption-based Grouper). As observed, our method frequently mis-assigns “Surfing” to the “Kayaking” cluster under the Activity criterion. Upon examining the intermediate criterion captions generated by the MLLM, we found that this error is largely due to the MLLM incorrectly describing a “Surfboard” as a “Kayak”. This highlights the importance of the MLLM’s ability to accurately describe images, as it is critical for the performance of our system. Potential improvements could include majority voting or model ensembling using different MLLM models.
Finally, we observed that our method sometimes fails to distinguish subtle, fine-grained differences between images, leading to incorrect labels. For example, as shown in Figure 16, “Edamame” or “Pho” are typical dishes from China, Vietnam, and Japan, but they may be presented differently depending on the cuisine. The “Edamame” shown in Figure 16 is presented in a traditional Japanese style, yet our model incorrectly predicted it as Chinese cuisine. This oversight of fine-grained details could be improved by employing a more advanced prompting strategy (Liu et al., 2024e).
Appendix J Further Study on Multi-granularity Clustering
In this section, we provide a detailed study on how different levels of multi-granularity output from our $\mathcal{X}$-Cluster framework impact grouping results. Specifically, for the Action-3c dataset, we employed human annotators to label two additional granularity levels for the criteria Action and Location. For the Action criterion, we consider the original annotation as fine-grained (L3) and tasked annotators to name the action in the image using more abstract and general coarse-grained (L1) and middle-grained (L2) labels. For the Location criterion, we consider the original annotation as middle-grained (L2) and tasked annotators to provide both more abstract coarse-grained (L1) labels and more specific fine-grained (L3) labels. This process resulted in expanded ground-truth annotations at three distinct semantic granularity levels for both the Action and Location criteria of the Action-3c dataset.
Figure 17: Further study on the influence of multi-granularity clustering output. We evaluate the CAcc and SAcc of the multi-granularity grouping results at each predicted clustering granularity level against each ground-truth annotation granularity level for the Action and Location criteria of the Action-3c dataset. The Harmonic Mean of CAcc and SAcc is reported for each granularity pair. L1, L2, and L3 represent the coarse-grained, middle-grained, and fine-grained levels, respectively, for both predictions and annotations.
Next, we quantitatively evaluated the multi-granularity grouping results at each predicted clustering granularity level against each ground-truth annotation granularity level by measuring clustering accuracy (CAcc) and semantic accuracy (SAcc). The main caption-based $\mathcal{X}$-Cluster framework was used for this experiment. In Figure 17, we report the Harmonic Mean of CAcc and SAcc for the Action and Location criteria of Action-3c, across each predicted clustering granularity level evaluated against each ground-truth annotation level. As clearly shown, the highest grouping performance consistently appears along the diagonal. This indicates that the best grouping performance is achieved when the predicted granularity matches the annotation granularity.
These experimental results not only highlight the importance of the multi-granularity output of our framework but also validate the effectiveness of our multi-granularity design in aligning with user-preferred granularities that is reflected by the annotations in these experiments.
Appendix K Study on Handling Invalid Criteria
At the criteria refinement step, invalid grouping criteria (False Positives) may be proposed due to hallucinations from large language models (LLMs). While we did not observe hallucinated criteria being introduced during our experiments across six datasets and three application studies, it is important to further investigate the potential impact of such invalid criteria on the proposed $\mathcal{X}$-Cluster system.
To this end, we design and conduct a control experiment using the Fruit-2c dataset (Muresan and Oltean, 2018), where we artificially introduced two “hallucinated” invalid grouping criteria (False Positives), Action and Clothing Style, into the refined criteria pool. These invalid criteria were then used in the subsequent grouping process to evaluate their effect on our system. We apply the main Caption-based Grouper to group fruit images based on these “hallucinated” criteria.
The grouping results for the two invalid criteria are presented in Tab. 31. As observed, when processing invalid “hallucinated” criteria, nearly all images are assigned to a cluster named “Not visible” by our framework. This occurs because, in the absence of relevant visual content in the images, the MLLM-generated captions do not include descriptors corresponding to the invalid criteria. Consequently, the LLM creates a “Not visible” cluster and assigns the images to it. Since the system provides interpretable outputs, users can easily identify and disregard such invalid groupings. This control experiment highlights the robustness of our system against hallucination in practical scenarios.
Table 31: Study of the Influence of Invalid Grouping Criteria (False Positives) on the Fruit-2c Dataset. We report the distribution of predicted groupings under the two “hallucinated” invalid grouping criteria. The main Caption-based Semantic Grouper is used for this experiment. †: For simplicity, all other minority clusters are grouped as “Others”.
Predicted Clusters Action (%) Clothing Style (%)
Not visible 98.3 96.7
Others† 1.7 3.3
Appendix L Study on Model Hallucination and Bias
Model hallucination. LLM hallucination (Wang et al., 2024c) typically occurs when LLMs are tasked with complex queries requiring world knowledge or factual information—for instance, answering a question like "Who was the 70th president of the United States?" might lead to a fabricated response. However, in our system, the use of LLMs is fully grounded in the visual descriptions (tags or captions) of the images. Consequently, the LLM output is strongly constrained to analyzing these visual descriptions, significantly reducing the likelihood of hallucination. That said, LLM hallucination can still have mild effects on clustering results. For example, as discussed in the failure case analysis in Appendix I, the LLM incorrectly grouped “Korean bibimbap” and “Vietnamese rice noodles” under “Chinese cuisine” (see Figure 16). MLLMs also play a crucial role in our system, as they are responsible for translating images into text for subsequent processing steps. MLLM hallucination (Wang et al., 2024c) typically involves incorrectly identifying the existence of objects, attributes, or spatial relationships within an image. However, since our proposed system operates at the dataset level rather than on a per-image basis, it is largely insensitive to such hallucinations, especially at the fine-grained visual detail level. Moreover, as our system is training-free, it can be further enhanced with LLM or MLLM hallucination mitigation techniques, such as the Visual Fact Checker (Ge et al., 2024), which we leave as a direction for future work.
Model Bias. Foundation models such as LLMs and MLLMs are known to inherit biases from their training data (Bommasani et al., 2021). In our system, we addressed potential biases using Hard Positive Prompting techniques: i) MLLM Bias Mitigation: The MLLM is further prompted to generate criterion-specific captions that focus solely on describing the criterion-related content in each image. This approach constrains the MLLM from generating irrelevant content influenced by inherent biases; ii) LLM Bias Mitigation: Similarly, when prompting the LLM to assign image captions to clusters, we condition it to concentrate exclusively on the Criterion depicted in each image (see Tab. 17).
To validate the effectiveness of these bias mitigation techniques, we conducted a fair clustering experiment. Specifically, following Kwon et al. (2024), we sampled images for four occupations (Craftsman, Laborer, Dancer, and Gardener) from the FACET (Gustafson et al., 2023) dataset, which contains images from 52 occupations. For each occupation, we selected 10 images of men and 10 images of women, totaling 80 images, ensuring a ground-truth gender proportion disparity of 0% for each occupation. Using our main $\mathcal{X}$-Cluster system, we grouped these images based on the criterion Occupation using three bias mitigation strategies: i) No mitigation: using general descriptions from the MLLM for LLM grouping; ii) Our default hard positive prompting strategy: using criterion-specific captions from the MLLM for LLM grouping; and iii) Our default strategy with additional negative prompt: adding a simple negative prompt, “Do not consider gender,” to both the MLLM captioning and LLM grouping prompts.
In this experiment, non-biased result is defined as achieving equal gender proportions within each cluster. Tab. 32 presents the average gender ratios of the clustering results for each method across the four occupations. As observed, without bias mitigation, $\mathcal{X}$-Cluster exhibits noticeable gender bias in the studied occupations, with a gender disparity of 19.4%. However, our default bias mitigation techniques effectively reduce this disparity to 4.9%, achieving performance comparable to the addition of a manual negative prompt. This experiment demonstrates the effectiveness of our bias mitigation strategy and highlights the potential for further reducing model bias in our framework using more advanced techniques.
Table 32: Average gender ratio and disparity across the four studied occupations (Craftsman, Laborer, Dancer, and Gardener) from the FACET dataset. Images sampled from each occupation have an equal proportion of genders. Results from different bias mitigation strategies are reported.
Bias Mitigation Strategy Male (%) Female (%) Gender Disparity (%)
Ground-truth 50.0 50.0 0.0
No mitigation 40.3 59.7 19.4
Ours (default) 47.6 52.5 4.9
Ours w. Negative prompt 48.4 51.6 3.2
Appendix M Computational Cost Analysis
The proposed $\mathcal{X}$-Cluster framework is training-free, requiring only inference processes. Specifically, our main framework (Caption-based) requires up to 31 GB of GPU memory to operate. All experiments reported in the paper were conducted on 4 Nvidia A100 40GB GPUs. In Tab. 33, we provide a detailed analysis of the computational efficiency of our main $\mathcal{X}$-Cluster framework (Caption-based Proposer and Caption-based Grouper) on the COCO-4c benchmark (5,000 images with four criteria) across various hardware configurations. For these experiments, we used LLaVA-NeXT-7B (Liu et al., 2024b) as the MLLM and Llama-3.1-8B (Meta, 2024a) as the LLM.
As shown in Tab. 33, organizing 5,000 images based on all four discovered criteria can be completed by $\mathcal{X}$-Cluster in 29.1 hours on a single A100 GPU or 16.7 hours on a single H100 GPU. More importantly, most steps in our framework, such as per-image captioning and per-caption cluster assignment, are parallelizable across multiple GPUs, significantly accelerating the process. Therefore, when parallelizing the framework on 4 A100 or H100 GPUs, we achieve approximately a 4$\times$ speedup, reducing computational time to 7.6 hours on 4 A100 GPUs and 4.3 hours on 4 H100 GPUs.
Table 33: Computational cost analysis on the COCO-4c benchmark (5,000 images with four criteria). We report the average and total time costs on various machines. The time costs were calculated for organizing all 5,000 images according to all the 4 criteria. Our main caption-based $\mathcal{X}$-Cluster framework is used in this experiment.
Method Hardware Average time cost (sec/img) $\downarrow$ Total time cost (hrs) $\downarrow$
$\mathcal{X}$-Cluster 1 Nvidia A100-40GB 20.9 29.1
4 Nvidia A100-40GB 5.5 7.6
1 Nvidia H100-80GB 12.0 16.7
4 Nvidia H100-80GB 3.1 4.3
Appendix N System Sensitivity Analysis of Various MLLMs and LLMs
@@ Appendix I Failure Case Analysis
Figure 18: Sensitivity analysis of different MLLMs and LLMs on the six OpenSMC benchmarks. Top (a): We fix the LLM to Llama-3.1-8B and study the impact of different MLLMs. Bottom (b): We fix the MLLM to LLaVA-NeXT-7B and study the impact of different LLMs. The average clustering accuracy(%) across different criteria is reported on the left, while the average semantic accuracy(%) is reported on the right.
@@
In Figure 18, we perform a system-level sensitivity analysis using our default system configuration (caption-based proposer and caption-based grouper) to examine the impact of different MLLMs and LLMs on the system performance. Since all variants successfully propose the basic criteria in each benchmark, we report the average clustering accuracy (CAcc) and semantic accuracy (SAcc) across various criteria for comparative analysis.
@@
Specifically, in Figure 18(a), we first fix the LLM in our system to Llama-3.1-8B (Meta, 2024a) and assess the influence of various MLLMs: GPT-4V (OpenAI, 2023), BLIP-3-4B (Xue et al., 2024), and LLaVA-NeXT-7B (Liu et al., 2024b). Next, in Figure 18(b), we set the MLLM to LLaVA-NeXT-7B and evaluate different LLMs: GPT-4-turbo (OpenAI, 2023), GPT-4o (OpenAI, 2024), Llama-3-8B (Meta, 2024b), and Llama-3.1-8B.
@@
Findings in Figure 18(a) indicate a direct correlation between the size of the MLLM and the ability of our system to uncover substructures, highlighting the significant role of MLLMs in translating visual information into natural language. On the other hand, this scalability demonstrates that our system can enhance performance with more robust MLLMs, thanks to its training-free design, which ensures compatibility with any MLLM. Despite this, we use LLaVA-NeXT-7B as our default MLLM due to its reproducibility, being open-source and unaffected by API changes, and its capacity for local deployment, which upholds privacy by not exposing sensitive image data to external entities.
@@
As for the LLMs, as depicted in Figure 18(b), despite GPT-4-turbo showing marginally superior performance, the open-source Llama-3.1-8B achieves similar results across benchmarks, making it our default LLM. Notably, except for the Card-2c dataset, system performance remains largely consistent regardless of the power of the LLM. This consistency suggests that the reasoning task for OpenSMC, given the capabilities of modern LLMs to tackle complex problems (Street et al., 2024), is relatively straightforward.
Appendix O Study on Fine-grained Image Collections
@@
Image collections may include fine-grained grouping criteria, such as Bird species in bird photography. Fine-grained criteria pose unique challenges for substructure discovery due to small inter-class differences and large intra-class variations (Zhang et al., 2014; Vedaldi et al., 2014; He and Peng, 2017). This requires the model to detect subtle visual distinctions to accurately infer cluster names and guide the grouping process. The straightforward captioning process in our current framework may not fully capture these subtle visual nuances. However, the modular design of our framework allows for seamless integration of advanced cross-modal chain-of-thought (CoT) prompting strategies to address this issue.
@@
Appendix P Further Details of the Application Study
@@
In this section, we present additional implementation details, evaluation results, and findings for the application study discussed in § 6 of the main paper. Specifically, § P.3 offers further evaluation results and implementation details on using our predicted distribution to train a debiased model with GroupDRO (Sagawa et al., 2020). § P.1 outlines the implementation of the user study that assesses the alignment between predicted biases and human judgments, along with comprehensive findings for all studied occupations and identified criteria. Finally, § P.2 provides additional insights from the analysis of social media image popularity.
P.1 Further Details on Discovering Novel Bias in Text-to-Image Diffusion Models
Image Generation for the Subject Occupation: Following prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016), we selected nine occupations for our study: three stereotypically biased towards females (Nurse, Cleaning staff, Call center employee), three biased towards males (CEO, Firefighter, Basketball player), and three considered gender-neutral (Teacher, Computer user, Marketing coordinator). We used two state-of-the-art T2I diffusion model, DALL·E3 (Betker et al., 2023) and Stable Diffusion (SDXL) (Podell et al., 2024) to generate 100 images for each occupation for our study. This resulted in a total of 1,800 images. For each occupation, we provide some examples of images generated by DALL·E3 in Figure 22, while provide some examples of images generated by SDXL in Figure 23. We only used the simple prompt “A portrait photo of a $<$OCCUPATION$>$” for image generation for all occupations and did not include any potential biases in the prompt.
where $\mathcal{D}_{\text{Occupation}}$ represents the generated images for each occupation, $\mathcal{C}^{l}$ denotes the clusters discovered under the $l$-th criterion, and $p(c^{l}|\mathcal{C}^{l})$ is the probability of each cluster under the current distribution. The resulting score is bounded between $\mathcal{H}_{bias}^{l}\in[0,1]$, where 0 indicates no bias towards a specific cluster (concept) under the evaluated criterion, and 1 indicates that the images are completely biased towards a particular cluster (concept) (e.g., “Grey” hair color) within the current bias dimension (e.g., Hair color). We used the score defined in 1 to quantify the biases for each occupation across the 10 discovered grouping criteria. We report the bias intensity score for each occupation and each model across the 10 discovered grouping criteria in Figure 19.
Human Evaluation Study Details: To assess the alignment between our method’s predictions and human judgments on bias detection, we conducted a user study to gather human evaluation results for the generated images. As shown in the questionnaire example in Figure 24, participants were presented with images generated by DALL-E3 and SDXL for each occupation and were asked to identify the bias direction (dominant class) for each of the 10 discovered criteria and rate the bias intensity on a scale from 0 to 10. We collected responses from 54 anonymous participants, resulting in 6 human evaluations for each occupation and each criterion.
The Absolute Mean Error (AME) between the bias intensity scores predicted by our system and those rated by humans (scaled to 0 to 1) was 0.1396. Additionally, our system’s predicted bias directions aligned with human evaluations 72.3% of the time, with most discrepancies occurring in the criteria of “Age group,” “Skin tone,” and “Accessories worn.” These findings indicate a strong correlation between our system’s predictions and human judgments, validating the effectiveness of our approach. Detailed user study results are provided in § P.1. We believe the discrepancies in certain criteria may be due to the influence of personal subjective cognition on respondents’ answers. In Figure 19, we present the human evaluation results, averaged across all participants for each model, occupation, and criterion, with the human ratings scaled from 0 to 1.
Complete Results and Additional Findings: In Figure 19, we present the detailed bias detection results for each model, occupation, and criterion, alongside human evaluation scores for reference. A particularly interesting phenomenon emerges: While DALL·E3 significantly outperforms SDXL on the well-known bias dimensions (e.g., Gender, Race, Age, and Skin tone), both DALL·E3 and SDXL exhibit moderate to strong biases along the novel bias dimensions (e.g., Hair color, Mood, Attire, and Accessories).
We speculate that DALL·E3’s superior performance in mitigating well-known biases may be attributed to its “guardrails” (OpenAI, 2022b), designed as part of its industrial deployment to avoid amplifying social biases via its easily accessible APIs. However, these guardrails do not prevent it from exhibiting biases along the novel dimensions discovered by our method, as these dimensions remain understudied. This observation highlights the importance of studying novel biases that could potentially exist in widely used T2I generative models to prevent further bias amplification.
P.2 Further Details on Analyzing Social Media Image Popularity
Figure 20: Complete analysis of social media photo popularity on the SPID dataset. We display the Top Trending and Top Mainstream clusters, along with the popularity distribution of data points within these clusters across all ten discovered criteria (in Grey).
With the rise of image-centric content on social media platforms like Instagram, Flickr, and TikTok, understanding what makes an image popular has become crucial for applications such as marketing, content curation, and recommendation systems. Traditional research often approaches image popularity as a regression problem (Ortis et al., 2019; Cheng et al., 2024), utilizing metadata like hashtags, titles, or follower counts. However, the specific semantic visual elements that contribute to an image’s popularity remain largely unexplored. In this study, we applied our proposed method to automatically categorize social media images based on semantic visual elements across different dimensions (criteria). By analyzing these interpretable results alongside image popularity metrics (e.g., number of views), we gained insights into the factors contributing to virality and identified common visual traits among popular images. These insights can provide valuable guidance for content creators and advertisers, enhancing productivity and informing strategic decision-making.
To expand on the discussion in § 6 of the main paper, we present the complete findings across all ten discovered criteria in Figure 20. Notably, we consistently observed a sharp semantic contrast between the visual elements in top trending images and those in the mainstream images across all ten criteria. For instance, there is a contrast between Urban sophisticated and Modern minimalist under Interior Design, Rustic architecture and Modern architecture under Architecture Style, and Event venues versus Urban residential areas under Location.
This recurring observation reinforces the idea that viral (or trending) content tends to capture more attention, likely because it features novel, surprising, or striking visual elements. Humans are inherently attracted to stimuli that deviate from the norm (Priester et al., 2004; Bruni et al., 2012; Palmer and Gore, 2014). On the other hand, widely uploaded yet “neutral” content is shared more often due to its familiarity and broad appeal, though it is less likely to provoke the strong emotional responses that fuel virality. We believe the insights generated by our method could offer valuable guidance to social media platform practitioners, helping them tailor their content more effectively to target audiences and gain a deeper understanding of social media image trends from various perspectives.
P.3 Confirming and Mitigating Dataset Bias
Confirming and Mitigating Dataset Bias: Given an image collection that contains spurious correlations (Geirhos et al., 2020), we are curious whether we can proactively find this issue caused by data bias directly from the training images without relying on either the annotations (Sagawa et al., 2020) or post hoc misclassified images (Kim et al., 2024). As a case study, we applied the proposed $\mathcal{X}$-Cluster framework to the 162k training images of the CelebA (Liu et al., 2015) dataset—a binary hair color classification dataset where the target label “Blond” is spuriously correlated with the demographic attribute “Female” in its training split.
Table 35: Debiasing Results and Comparison on CelebA. We use the groups discovered by $\mathcal{X}$-Cluster to train DRO and compare it with state-of-the-art debiasing methods. Additionally, we present DRO results using the ground-truth distribution (DRO+GT) for reference.
Method Worst Avg.
JTT 81.5 88.1
CNC 88.8 89.9
DRO+B2T 90.4 93.2
DRO+$\mathcal{X}$-Cluster 90.9 93.1
DRO+GT 89.7 93.6
Findings: As expected, our method identified the grouping criteria Hair color and Gender. Next, we analyzed the predicted gender distributions within the “Blond” and “Not Blond” (all other colors) clusters. As shown in Figure 21, we observed that the gender distribution within the “Blond” cluster is highly skewed, with 86.5% of the images representing females, closely matching the ground-truth distribution (94.3%). Such an imbalance confirms the potential issue of spurious correlations between “Blond” and “Female”. To further validate this observation, following B2T (Kim et al., 2024), we used our predicted distributions to train a debiased model with GroupDRO (Sagawa et al., 2020) and compared it with other unsupervised bias mitigation methods, including JTT (Liu et al., 2021), CNC (Zhang et al., 2022), B2T, and GroupDRO trained with ground-truth labels. As shown in Tab. 35, our debiased model achieved robust performance, comparable to that of B2T, demonstrating the reliability of its discovered distributions.
@@
where $H(A)$ and $H(Y)$ represent the normalized entropy of the gender and hair color distributions, respectively. A value of $H(A)$ or $H(Y)$ equal to 1 indicates a uniform distribution (i.e., no class imbalance). We then used the ground-truth distribution from the dataset’s labels and our predicted distribution to estimate the spurious correlation intensity using the score from 2. For gender and hair color, our method’s predictions yielded a score of $I_{Pred}=0.10$, which is nearly identical to the ground-truth score of $I_{GT}=0.11$. This demonstrates that our method effectively identifies and confirms the bias directly from the training set.